Weather records show that the probability of rainfall in Florida in June is 0.86. What is the probability that it rains there in June at least 9 times in a decade?

Verified Answer :

R - it is rain  (probablity: 0.86)
N - it is not rain (probablity: 1-0.86 = 0.14)
You've got these cases:

R R R R R R R R R N
R R R R R R R R N R
R R R R R R R N R R 
. . . 
N R R R R R R R R R     -  it is 10 cases.  Probablity each case is
(0.86)^9 * 0.14 - it is approximately  0.036   (9R and 1N - you multiply everything)

It is 10 cases so you have to multiply by 10  and you get 0.36

There is also eleventh case:

R R R R R R R R R R 

Probablity is (0.86)^10 approximately 0.22

Add these cases and you'll get your probablity:  0.36 + 0.22 = 0.58

0.5816 = 58.16% probability that it rains there in June at least 9 times in a decade.

For each June 9 in Florida, there are only two possible outcomes. Either it rains, or it does not. The probability of raining on June 9 of an year is independent of rain on June 9 on any other year, which means that the binomial probability distribution is used to solve this question.

---------------------------------------

Binomial probability distribution

The binomial probability is the probability of exactly x successes on n repeated trials, and X can only have two outcomes.

\(P(X = x) = C_{n,x}.p^{x}.(1-p)^{n-x}\)

In which \(C_{n,x}\) is the number of different combinations of x objects from a set of n elements, given by the following formula.

\(C_{n,x} = \frac{n!}{x!(n-x)!}\)

And p is the probability of X happening.

---------------------------------------

  • 0.86 probability of rain means that \(p = 0.86\)
  • A decade means 10 years, so \(n = 10\)

---------------------------------------

What is the probability that it rains there in June at least 9 times in a decade?

This is:

\(P(X \geq 9) = P(X = 9) + P(X = 10)\)

In which

\(P(X = x) = C_{n,x}.p^{x}.(1-p)^{n-x}\)

\(P(X = 9) = C_{10,9}.(0.86)^{9}.(0.14)^{1} = 0.3603\)

\(P(X = 10) = C_{10,10}.(0.86)^{10}.(0.14)^{0} = 0.2213\)

Thus

\(P(X \geq 9) = P(X = 9) + P(X = 10) = 0.3603 + 0.2213 = 0.5816\)

0.5816 = 58.16% probability that it rains there in June at least 9 times in a decade.

A similar question is found at https://brainly.com/question/9631195